If the larger of two consecutive even numbers is subtracted from twice the larger, the result is 12. Find the two numbers.
Answer by RAY100(1637)

Let x = 1st even no
let (x+2) = 2nd even no
2L-L=12
2(x+2)-(x+2) =12
(x+2)=12
subst 2 each side
x=10
check 10 is smaller no, 12 is larger no
2*12-12=12 ok